# Interplay between structural, electronic and topological properties in low-dimensional tellurium

### v2

@@ 2 Computational details
Our investigation employs first-principles calculations performed with the Vienna Ab initio Simulation Package (VASP)1, 2. The exchange–correlation potential is treated using both the generalized gradient approximation (GGA), parameterized by Perdew, Burke, and Ernzerhof (PBE)3, and the hybrid Heyd–Scuseria–Ernzerhof (HSE06) functional.4 Long-range van der Waals (vdW) interactions are included via the DFT-D3 correction method proposed by Grimme.5 Core–valence electron interactions are described using the projector augmented-wave (PAW) method,6 and the Kohn–Sham single-electron wavefunctions are expanded in a plane-wave basis set with a kinetic energy cutoff of 520 eV.
All atomic structures are fully relaxed until the residual forces on each atom are smaller than $1\times 10^{-6}$ eV/Å. To avoid spurious interactions arising from periodic boundary conditions, a vacuum spacing of 12 Å is introduced—along the $z$-direction for monolayers and along the transverse ($x$ and $y$) directions for nanowires. The Brillouin zone is sampled using k-point meshes generated by the Monkhorst–Pack scheme,7 with grids of ($15\times 15\times 8$) for bulk Te, ($13\times 17\times 1$) for $\alpha$-tellurene, ($9\times 12\times 1$) for $\beta$-tellurene, ($13\times 17\times 1$) for hexagonal planar, ($24\times 24\times 1$) for hexagonal buckled, ($13\times 13\times 1$) for pentagonal, ($17\times 17\times 1$) for Lieb-like, ($21\times 21\times 1$) for planar kagome, ($19\times 19\times 1$) for buckled kagome, ($23\times 23\times 1$) for square planar, ($23\times 23\times 1$) for buckled square, and ($1\times 1\times 18$) for nanowires.
Vibrational and thermodynamic properties are evaluated from phonon dispersion calculations using the finite-displacement method. Supercells of sizes ($4\times 4\times 4$) for bulk, ($4\times 4\times 1$) for monolayers, and ($1\times 1\times 15$) for nanowires are employed. The same plane-wave cutoff energy (520 eV) and k-point densities as in the electronic structure calculations are used, with a force convergence threshold of $10^{-6}$ eV/Å. Phonon dispersion relations and corresponding density of states (DOS) are computed using the Phonopy package.8, 9
To evaluate the Chern number and the $\mathbb{Z}_{2}$ topological invariant, maximally localized Wannier functions (MLWFs) are constructed using the Wannier90 package.10 The corresponding topological invariants are subsequently computed with the WannierTools post-processing code.11 The topological properties of the one-dimensional nanowire were analyzed using the modern theory of polarization. Atomic structures are visualized using the VESTA software,12 and all plots and graphical analyses are produced with the Matplotlib library.13
@@ 3 Results and Discussions
Tellurium crystallizes in a stable trigonal phase composed of one-dimensional helical chains of Te atoms along the $c$-axis. Each atom has a $5s^{2}5p^{4}$ configuration, where two $5p$ electrons form covalent bonds within the chains, while the $5s$ electrons remain core-like. The remaining $5p$ electrons form lone pairs oriented between chains, leading to interchain vdW interactions. This results in strong intrachain bonding and a quasi-layered, anisotropic structure 14, 15.
@@
The helical structure exhibits structural chirality, adopting either the right-handed $P3_{1}21$ ($D_{3}^{4}$) or the left-handed $P3_{2}21$ ($D_{3}^{6}$) space group, making them non-centrosymmetric systems. After structure optimization shown in Fig. 1, the obtained lattice parameters within GGA-PBE are $a=b=4.41$ Å and $c=5.93$ Å, with a Te-Te bond length of $d_{h}=2.90$ Å. These values are very close to previous theoretical calculations 16, 17, 18 and close to the experimental values of $a=b=4.45$ Å and $c=5.93$ Å 19.
The dynamic stability of trigonal tellurium (Te-I) is investigated through phonon dispersion calculations within the harmonic approximation, employing the finite displacement method. The calculated phonon band structure of trigonal tellurium (Te-I) along with thermal properties are shown in Fig. S1A closer inspection of the optical phonon modes at the $\Gamma$ point reveals doubly degenerate branches along the $\Gamma$-A direction at approximately 2.44 THz and 3.89 THz. These modes involve coupled atomic displacements combining bond stretching and angular distortions, and reflect the intrinsic chiral symmetry of the trigonal structure. While such degeneracies are compatible with chiral lattice dynamics, we emphasize that a definitive identification of Weyl phonons 20 would require an explicit evaluation of phonon Berry curvature and topological charges, which is beyond the scope of the present work. The observed phonon features therefore suggest, but do not by themselves establish, topologically nontrivial phononic behavior.
Much less explored are the 2D allotropes of tellurium predominantly found in the trigonal $\alpha$-phase ($\alpha$-Te), characterized by the $P\bar{3}m1$ space group, and the monoclinic $\beta$-phase ($\beta$-Te), belonging to the $P2/m$ space group with distinct zigzag and armchair directions. Tellurium propensity to form these 2D monolayers is attributed to its outer valence electron configuration and a potential Peierls instability, a distortion of the periodic lattice in a one-dimensional crystal that breaks its perfect translational symmetry, in this case, tellurium helical chains 21, 22. This instability can drive a spontaneous structural transition toward energetically more favorable 2D configurations, resulting in the formation of $\alpha$-phase, show in Fig. 2(a). This structure exhibits a lower total energy, than the $\beta$-phase shown in Fig. 2(b). Additionally, we investigate the following 2D-tellurium structures: Fig. 2(c) (buckled pentagonal), Fig. 2(d) (buckled kagome), Fig. 2(e) (buckled square), Fig. 2(f) (planar hexagonal), Fig. 2(g) (Lieb-like), Fig. 2(h) (planar kagome) and Fig. 2(i) (planar square). These forms of tellurium are proposed here inspired by novel newly synthesized 2D square/rectangular 23 and hexagonal 24 phases, kagome metals 25, 26, pentagonal bismuthene 27, graphene28, 29 and phosphorene 30 just to cite a few.
@@
For $\alpha$-Te, the calculated lattice parameters are $a=b=4.22\,\text{\AA }$, with a Te-Te bond length of $d=3.03\,\text{\AA }$, very similar to previous theoretical 16, 31, 32 and experimental 22, 33values. The $\beta$-tellurene phase exhibits lattice parameters $a=5.61\,\text{\AA }$ and $b=4.22\,\text{\AA }$, featuring Te-Te bond lengths of $d=3.03\,\text{\AA }$ and $d=2.76\,\text{\AA }$, similar to previous theoretical 34, 16, 32 and experimental data 35. The buckled pentagonal structure has lattice parameters of 7.71Å, the buckled kagome of 5.51Å and the buckled square of 4.10Å. This buckling leads to Te-Te bond lengths of 3.02 Å, 2.96Å and 3.03Å, respectively.
Topology is conditional on stabilization. The buckled kagome and buckled square tellurene lattices are found to host nontrivial $Z_{2}$ = 1 topology arising from SOC-gapped near-crossings in the electronic structure. Phonon calculations indicate that these free-standing configurations exhibit soft modes, suggesting that their stabilization likely requires interaction with a substrate or external constraints. Importantly, similar kagome- and square-based tellurium phases have been experimentally realized on metallic substrates, supporting the physical relevance of these structures. Our results therefore establish the intrinsic topological character of these lattices, conditional on structural stabilization. Applying 5% strain modifies the band dispersions and shifts the band edges, but SOC continues to gap the near-crossings along $\Gamma$-X-M-$\Gamma$ [Fig. S6(l)]. For the buckled structures, our results are comparable to previous ones for $\alpha$-Te, $\beta$-Te 36, pentagonal 37, hexagonal 24. We emphasize the importance of the substrate: the DFT ground state is a square lattice, while in the experiment a rectangular reconstruction is found on Ni(111) substrate 23.
@@
In $\alpha$-Te, one of the acoustic branches exhibits a soft mode [Fig. S2(a)], indicating a possible dynamic instability under small perturbations. In $\beta$-Te, all three acoustic modes are softened, particularly along the armchair direction ($\Gamma$–X), and several optical branches show a noticeable reduction in energy [Fig. S2(b)]. This softening implies a higher susceptibility of these 2D structures to structural distortions or phase transitions when subjected to external perturbations. Although the buckled pentagonal, [Fig. S2(c)], and buckled kagome phases show small imaginary frequencies, as seen in Fig. SS2(d) and (e), it is be possible to stabilize these phases on an appropriate substrate. Strain induced by the substrate possibly introduce some electrostatic interaction and change the corrugation. This is consistent with experimental results for pentagonal37, hexagonal,24, and rectangular lattices23 which are reported to be topological systems.
@@
We now turn our discussion to single-helix tellurium nanowires (Te-h). These one-dimensional structures can be obtained, for example, by decoupling the helical atomic chains that constitute bulk trigonal tellurium (Te-I),38, 39 thereby preserving the intrinsic geometrical chirality and screw symmetry of Te-I at the single-chain level while removing the interchain packing present in the bulk crystal. The resulting quantum-confined geometry, combined with the strong spin–orbit coupling of tellurium, gives rise to distinctive one-dimensional electronic properties. We emphasize that the structural chirality of the helical chain should not be confused with chiral (sublattice) symmetry of the electronic Hamiltonian; in the present nanowires the latter is absent (see Fig. S7 in SI and the details for Zak phase calculations.
@@
Fig. 3 shows the optimized structures of an ultra-thin tellurium nanowire. The obtained lattice parameter is $c=5.67\,\text{\AA }$ with a Te-Te bond length of $d=2.74$ Å  in good agreement with previous theoretical results 40, 18. From the phonon dispersion curves depicted in Fig. S5, the absence of imaginary frequencies in Te-h along the high symmetry paths, indicating the possibility of local stability at low temperature, in agreement with previously reported theoretical data 40. Also, we see that the acoustic branches and one optical branch have softened, since their energies decrease compared to Te-I ($\Gamma$ - A), suggesting potential structural phase transitions under small perturbations. On the other hand, the energy of some optical branches increased significantly (hardening). For example, the mode at 0.14 THz [Fig. S5(a)] corresponds to torsional oscillations, similarly to the Te-I phase, but with a significantly lower eigenfrequency. This reduction is attributed to the absence of inter-chain interactions in the system. The Te-h phase hardens the modes that involve a combination of bond stretching and bending. The shorter Te–Te bond length in Te-h (2.74 Å) as opposed to bulk Te (2.90 Å) is consistent with this behavior.
@@
According to Table 1, $\alpha$-Te is identified as the most stable 2D phase, while the remaining phases are metastable with respect to $\alpha$-Te. The smaller cohesive energy of $\alpha$-Te compared to $\beta$-Te confirms its higher thermodynamic stability. Although $\beta$-Te is thermodynamically less stable than $\alpha$-Te, the monoclinic structure may still be experimentally accessible under specific synthesis conditions,41. The buckled pentagonal, kagome, and square structures exhibit comparable cohesive energies, in agreement with experimental observations for these phases obtained under different substrates and growth conditions,37, 24, 23. This finding suggests that alternative tellurium phases can be stabilized depending on the experimental environment and synthesis parameters.
@@
The MLWF-HSE06 band structures calculated with and without spin–orbit coupling (SOC), shown in Fig. SS6(j) for Te-I, reveal an indirect electronic band gap located at the high-symmetry point H. The computed band gaps are 0.49 (0.30) eV with (without) SOC, in good agreement with the experimental value of 0.33 eV,45. These results classify Te-I as a narrow-gap semiconductor. Within the energy range from 1-1 eV to 0 eV, the valence band at the H point is fourfold degenerate when SOC is neglected. The inclusion of SOC lifts this degeneracy, resulting in two non-degenerate states and one doubly degenerate state at lower energy. The region between the high-symmetry points L–H–A highlights features of the conduction band where doubly degenerate states exhibit linear dispersion—indicative of Weyl nodes located close to the Fermi level.
@@
Figs. S6(a) and (b) present the electronic band structures of $\alpha$-Te and $\beta$-Te, respectively, calculated using the MLWF-HSE06 exchange–correlation functional, both with and without spin–orbit coupling (SOC). The inclusion of SOC leads to a reduction in the band gap for both phases. Specifically, for $\alpha$-Te, the band gap decreases from $1.04$ eV to $0.75$ eV upon inclusion of SOC, in excellent agreement with previous theoretical reports,46, 47, 48. Similarly, for $\beta$-Te, the band gap decreases from $1.77$ eV to $1.44$ eV with SOC, consistent with earlier theoretical studies,46, 48, 49, 50.
@@
Nevertheless, it is important to note that non-trivial topological phases can arise under suitable external conditions. Topological phase transitions may be induced by perturbations such as mechanical strain, magnetic impurities, or doping. As discussed in Ref. 51, for instance, applying isotropic strain can drive tellurene from a trivial state into a topological phase. For completeness, the $\langle S_{x}\rangle$ spin components of the: pentagonal phase are shown in Figs. 7(c,d), buckled kagome phase [Figs. 7(e,f)], and buckled square phase [Figs. 7(g,h)].
@@
Fig. 8(a) presents the band structure of the Te helicoidal nanowire (Te-h), computed using the MLWF-HSE06 functional both with and without SOC. The inclusion of SOC consistently narrows the band gap, reducing it from 2.49 eV to 2.23 eV. These values align well with previous theoretical reports,18. In both cases, the band structure displays predominantly flat dispersion. This combination of a finite band gap and quasi-flat bands makes Te-h an appealing material for photonic applications,52. Fig. 8(b) shows the inclusion of SOC lifts band degeneracies in a manner similar to that observed in Te-I along the $\Gamma$–A direction. Along the high-symmetry path of Te-h, four band crossings are identified: two occurring at the same $K$-point—P1 in the valence band at $2.7,\text{eV}$ and P2 in the conduction band near $2,\text{eV}$—and two additional crossings near the $\Gamma$-point, P3 at $2.7,\text{eV}$ and P4 around $2,\text{eV}$. Fig. 8(c) presents the PDOS for Te-h, showing that the $p_{x}$, $p_{y}$, and $p_{z}$ orbitals contribute comparably to both the VBM and CBM. However, the CBM exhibits a more pronounced contribution from the $p_{x}$ orbital. Due to quantum confinement, Te-h displays highly localized electronic states, which is reflected in the discrete features of the band structure and in the pronounced peaks observed in the DOS.
@@
The effective masses [Table S2] were extracted from the band structure near the inflection points at the VBM and CBM. Te-h effective masses are suggestive of potentially high mobility53, 54, 55. Moreover, the modest reduction in electron and hole mobility in Te-h can help mitigate current leakage in nanoelectronic devices 56. In units of the free-electron mass, the electron effective mass of Te-h is 0.484, while the hole effective mass is 0.817 (see Table S2).
In Table 2, the topological invariant $\mathbb{Z}_{2}$ for the 2D phases of tellurium is presented. The invariants are calculated following the method described in Ref.,11. The buckled kagome and buckled square phases exhibit non-trivial topology, whereas $\alpha$-Te, $\beta$-Te, and the buckled pentagonal phase are topologically trivial. These topological classifications are consistent with the spin textures shown in Fig. 7.
Using experimentally reported lattice parameters24, the planar hexagonal tellurene phase is found to exhibit semimetallic behavior and a nontrivial $\mathbb{Z}_{2}=1$ index when SOC is included, consistent with previous experimental observations. However, our phonon calculations indicate that the free-standing planar hexagonal lattice is dynamically unstable. We therefore investigated stabilization mechanisms through strain engineering and surface functionalization. Under 5% isotropic in-plane strain, as well as under one-side hydrogen passivation, the system undergoes a transition to a gapped electronic state while preserving a nontrivial $\mathbb{Z}_{2}=1$ invariant. In these stabilized configurations, the SOC-induced gap and the winding of the Wilson loop unambiguously identify a quantum spin Hall phase.
@@
For the 2D phases, the effective masses for electrons (holes) in $\alpha$-Te are 0.108 (0.135). In $\beta$-Te, the electron effective masses are 1.009 (along X) and 0.203 (along Y), while the hole effective masses are 0.368 (X) and 0.127 (Y), indicating strong transport anisotropy. Our results suggest that both $\alpha$-Te and $\beta$-Te could have higher electron and hole mobilities than structurally or symmetrically similar materials such as 2H-MoS2 57, 58 and phosphorene 59, 60. While $\alpha$-Te exhibits higher mobilities for both charge carriers overall, $\beta$-Te shows pronounced anisotropy due to its geometry, with significantly reduced mobility along the armchair direction ($\Gamma$–X) compared to the zigzag direction ($\Gamma$–Y).

### v3

@@ 1 Introduction
Tellurium (Te) is a rare, silvery metalloid found only in trace quantities in Earth’s crust and seawater. Despite its scarcity, Te has attracted increasing attention owing to its distinctive physical and chemical properties, particularly as a semiconductor with emerging technological relevance. As a group-16 chalcogen, Te exhibits remarkable crystalline and electronic characteristics, including non-trivial topological behavior driven by strong spin–orbit coupling (SOC).45, 70
In its bulk phase, Te crystallizes in a trigonal structure composed of one-dimensional (1D) chiral helical chains aligned along the $c$-axis. This non-centrosymmetric motif breaks inversion symmetry and gives rise to a rich landscape of topological phenomena. Trigonal tellurium is a narrow-gap semiconductor widely recognized as hosting Weyl nodes in its bulk band structure.22, 38, 46 The characteristic hedgehog spin texture in momentum space further emphasizes the critical role of inversion-symmetry breaking and SOC in shaping Te electronic topology.70
[…]
Tellurium can also form one-dimensional nanostructures such as nanowires and nanoribbons. First synthesized via hydrothermal methods,39 these systems preserve the helical atomic backbone of bulk Te, offering an ideal platform to probe chirality-driven physics under quantum confinement. The recent fabrication of ultrathin, high-crystallinity Te nanowires 13 underscores the growing interest in identifying low-dimensional electronic and boundary-state phenomena in 1D Te architectures.
The interplay of structural flexibility, chirality, strong SOC, and symmetry breaking thus positions tellurium as a uniquely versatile platform for realizing and engineering emergent topological phases.71, 28
Despite these advances, a unified understanding of how topology evolves across Te dimensionalities remains incomplete. Here, we address this gap through first-principles calculations of tellurium in its 3D, 2D, and 1D forms, elucidating how chirality, symmetry breaking, and SOC govern its electronic topology. We confirm the existence of Weyl nodes in bulk Te, reveal that 2D allotropes exhibit symmetry-tunable electronic structures, and demonstrate that helical Te nanowires retain bulk chirality while hosting edge-localized states associated with quantum confinement.
This work establishes symmetry-driven connections between the Weyl physics of bulk Te and its low-dimensional derivatives. Beyond advancing the fundamental understanding of symmetry-protected states in reduced dimensions, our findings highlight tellurium as a promising candidate for next-generation spintronic, optoelectronic, and quantum devices, where topological phases can be engineered via strain, confinement, and chemical modification.
@@ 2 Computational details
Our investigation employs first-principles calculations performed with the Vienna Ab initio Simulation Package (VASP)37, 36. The exchange–correlation potential is treated using both the generalized gradient approximation (GGA), parameterized by Perdew, Burke, and Ernzerhof (PBE)51, and the hybrid Heyd–Scuseria–Ernzerhof (HSE06) functional.20 Long-range van der Waals (vdW) interactions are included via the DFT-D3 correction method proposed by Grimme.18 Core–valence electron interactions are described using the projector augmented-wave (PAW) method,7 and the Kohn–Sham single-electron wavefunctions are expanded in a plane-wave basis set with a kinetic energy cutoff of 520 eV.
All atomic structures are fully relaxed until the residual forces on each atom are smaller than $1\times 10^{-6}$ eV/Å. To avoid spurious interactions arising from periodic boundary conditions, a vacuum spacing of 12 Å is introduced—along the $z$-direction for monolayers and along the transverse ($x$ and $y$) directions for nanowires. The Brillouin zone is sampled using k-point meshes generated by the Monkhorst–Pack scheme,48 with grids of ($15\times 15\times 8$) for bulk Te, ($13\times 17\times 1$) for $\alpha$-tellurene, ($9\times 12\times 1$) for $\beta$-tellurene, ($13\times 17\times 1$) for hexagonal planar, ($24\times 24\times 1$) for hexagonal buckled, ($13\times 13\times 1$) for pentagonal, ($17\times 17\times 1$) for Lieb-like, ($21\times 21\times 1$) for planar kagome, ($19\times 19\times 1$) for buckled kagome, ($23\times 23\times 1$) for square planar, ($23\times 23\times 1$) for buckled square, and ($1\times 1\times 18$) for nanowires.
Vibrational and thermodynamic properties are evaluated from phonon dispersion calculations using the finite-displacement method. Supercells of sizes ($4\times 4\times 4$) for bulk, ($4\times 4\times 1$) for monolayers, and ($1\times 1\times 15$) for nanowires are employed. The same plane-wave cutoff energy (520 eV) and k-point densities as in the electronic structure calculations are used, with a force convergence threshold of $10^{-6}$ eV/Å. Phonon dispersion relations and corresponding density of states (DOS) are computed using the Phonopy package.62, 63
To evaluate the Chern number and the $\mathbb{Z}_{2}$ topological invariant, maximally localized Wannier functions (MLWFs) are constructed using the Wannier90 package.52 The corresponding topological invariants are subsequently computed with the WannierTools post-processing code.64 The topological properties of the one-dimensional nanowire were analyzed using the modern theory of polarization. Atomic structures are visualized using the VESTA software,47 and all plots and graphical analyses are produced with the Matplotlib library.27
@@ 3 Results and Discussions
Tellurium crystallizes in a stable trigonal phase composed of one-dimensional helical chains of Te atoms along the $c$-axis. Each atom has a $5s^{2}5p^{4}$ configuration, where two $5p$ electrons form covalent bonds within the chains, while the $5s$ electrons remain core-like. The remaining $5p$ electrons form lone pairs oriented between chains, leading to interchain vdW interactions. This results in strong intrachain bonding and a quasi-layered, anisotropic structure 14, 33.
@@
The helical structure exhibits structural chirality, adopting either the right-handed $P3_{1}21$ ($D_{3}^{4}$) or the left-handed $P3_{2}21$ ($D_{3}^{6}$) space group, making them non-centrosymmetric systems. After structure optimization shown in Fig. 1, the obtained lattice parameters within GGA-PBE are $a=b=4.41$ Å and $c=5.93$ Å, with a Te-Te bond length of $d_{h}=2.90$ Å. These values are very close to previous theoretical calculations 77, 65, 35 and close to the experimental values of $a=b=4.45$ Å and $c=5.93$ Å 29.
The dynamic stability of trigonal tellurium (Te-I) is investigated through phonon dispersion calculations within the harmonic approximation, employing the finite displacement method. The calculated phonon band structure of trigonal tellurium (Te-I) along with thermal properties are shown in Fig. S1. A closer inspection of the optical phonon modes at the $\Gamma$ point reveals doubly degenerate branches along the $\Gamma$-A direction at approximately 2.44 THz and 3.89 THz. These modes involve coupled atomic displacements combining bond stretching and angular distortions, and reflect the intrinsic chiral symmetry of the trigonal structure. While such degeneracies are compatible with chiral lattice dynamics, we emphasize that a definitive identification of Weyl phonons 74 would require an explicit evaluation of phonon Berry curvature and topological charges, which is beyond the scope of the present work. The observed phonon features therefore suggest, but do not by themselves establish, topologically nontrivial phononic behavior.
Much less explored are the 2D allotropes of tellurium predominantly found in the trigonal $\alpha$-phase ($\alpha$-Te), characterized by the $P\bar{3}m1$ space group, and the monoclinic $\beta$-phase ($\beta$-Te), belonging to the $P2/m$ space group with distinct zigzag and armchair directions. Tellurium propensity to form these 2D monolayers is attributed to its outer valence electron configuration and a potential Peierls instability, a distortion of the periodic lattice in a one-dimensional crystal that breaks its perfect translational symmetry, in this case, tellurium helical chains 60, 25. This instability can drive a spontaneous structural transition toward energetically more favorable 2D configurations, resulting in the formation of $\alpha$-phase, show in Fig. 2(a). This structure exhibits a lower total energy, than the $\beta$-phase shown in Fig. 2(b). Additionally, we investigate the following 2D-tellurium structures: Fig. 2(c) (buckled pentagonal), Fig. 2(d) (buckled kagome), Fig. 2(e) (buckled square), Fig. 2(f) (planar hexagonal), Fig. 2(g) (Lieb-like), Fig. 2(h) (planar kagome) and Fig. 2(i) (planar square). These forms of tellurium are proposed here inspired by novel newly synthesized 2D square/rectangular 53 and hexagonal 40 phases, kagome metals S. Chaudhary, A. Consiglio, J. Singh, D. Di Sante, R. Thomale, and Y. Singh (2025), 21, pentagonal bismuthene 42, graphene72, 16 and phosphorene 61 just to cite a few.
@@
For $\alpha$-Te, the calculated lattice parameters are $a=b=4.22\,\text{\AA }$, with a Te-Te bond length of $d=3.03\,\text{\AA }$, very similar to previous theoretical 77, 76, 15 and experimental 25, 8values. The $\beta$-tellurene phase exhibits lattice parameters $a=5.61\,\text{\AA }$ and $b=4.22\,\text{\AA }$, featuring Te-Te bond lengths of $d=3.03\,\text{\AA }$ and $d=2.76\,\text{\AA }$, similar to previous theoretical 58, 77, 15 and experimental data 12. The buckled pentagonal structure has lattice parameters of 7.71Å, the buckled kagome of 5.51Å and the buckled square of 4.10Å. This buckling leads to Te-Te bond lengths of 3.02 Å, 2.96Å and 3.03Å, respectively.
Topology is conditional on stabilization. The buckled kagome and buckled square tellurene lattices are found to host nontrivial $Z_{2}$ = 1 topology arising from SOC-gapped near-crossings in the electronic structure. Phonon calculations indicate that these free-standing configurations exhibit soft modes, suggesting that their stabilization likely requires interaction with a substrate or external constraints. Importantly, similar kagome- and square-based tellurium phases have been experimentally realized on metallic substrates, supporting the physical relevance of these structures. Our results therefore establish the intrinsic topological character of these lattices, conditional on structural stabilization. Applying 5% strain modifies the band dispersions and shifts the band edges, but SOC continues to gap the near-crossings along $\Gamma$-X-M-$\Gamma$ [Fig. S6(l)]. For the buckled structures, our results are comparable to previous ones for $\alpha$-Te, $\beta$-Te 43, pentagonal 73, hexagonal 40. We emphasize the importance of the substrate: the DFT ground state is a square lattice, while in the experiment a rectangular reconstruction is found on Ni(111) substrate 53.
@@
In $\alpha$-Te, one of the acoustic branches exhibits a soft mode [Fig. S2(a)], indicating a possible dynamic instability under small perturbations. In $\beta$-Te, all three acoustic modes are softened, particularly along the armchair direction ($\Gamma$–X), and several optical branches show a noticeable reduction in energy [Fig. S2(b)]. This softening implies a higher susceptibility of these 2D structures to structural distortions or phase transitions when subjected to external perturbations. Although the buckled pentagonal, [Fig. S2(c)], and buckled kagome phases show small imaginary frequencies, as seen in Fig. SS2(d) and (e), it is be possible to stabilize these phases on an appropriate substrate. Strain induced by the substrate possibly introduce some electrostatic interaction and change the corrugation. This is consistent with experimental results for pentagonal73, hexagonal,40, and rectangular lattices53 which are reported to be topological systems.
@@
We now turn our discussion to single-helix tellurium nanowires (Te-h). These one-dimensional structures can be obtained, for example, by decoupling the helical atomic chains that constitute bulk trigonal tellurium (Te-I),49, 10 thereby preserving the intrinsic geometrical chirality and screw symmetry of Te-I at the single-chain level while removing the interchain packing present in the bulk crystal. The resulting quantum-confined geometry, combined with the strong spin–orbit coupling of tellurium, gives rise to distinctive one-dimensional electronic properties. We emphasize that the structural chirality of the helical chain should not be confused with chiral (sublattice) symmetry of the electronic Hamiltonian; in the present nanowires the latter is absent (see Fig. S7 in SI and the details for Zak phase calculations.
@@
Fig. 3 shows the optimized structures of an ultra-thin tellurium nanowire. The obtained lattice parameter is $c=5.67\,\text{\AA }$ with a Te-Te bond length of $d=2.74$ Å  in good agreement with previous theoretical results 17, 35. From the phonon dispersion curves depicted in Fig. S5, the absence of imaginary frequencies in Te-h along the high symmetry paths, indicating the possibility of local stability at low temperature, in agreement with previously reported theoretical data 17. Also, we see that the acoustic branches and one optical branch have softened, since their energies decrease compared to Te-I ($\Gamma$ - A), suggesting potential structural phase transitions under small perturbations. On the other hand, the energy of some optical branches increased significantly (hardening). For example, the mode at 0.14 THz [Fig. S5(a)] corresponds to torsional oscillations, similarly to the Te-I phase, but with a significantly lower eigenfrequency. This reduction is attributed to the absence of inter-chain interactions in the system. The Te-h phase hardens the modes that involve a combination of bond stretching and bending. The shorter Te–Te bond length in Te-h (2.74 Å) as opposed to bulk Te (2.90 Å) is consistent with this behavior.
@@
According to Table 1, $\alpha$-Te is identified as the most stable 2D phase, while the remaining phases are metastable with respect to $\alpha$-Te. The smaller cohesive energy of $\alpha$-Te compared to $\beta$-Te confirms its higher thermodynamic stability. Although $\beta$-Te is thermodynamically less stable than $\alpha$-Te, the monoclinic structure may still be experimentally accessible under specific synthesis conditions,23. The buckled pentagonal, kagome, and square structures exhibit comparable cohesive energies, in agreement with experimental observations for these phases obtained under different substrates and growth conditions,73, 40, 53. This finding suggests that alternative tellurium phases can be stabilized depending on the experimental environment and synthesis parameters.
@@
The MLWF-HSE06 band structures calculated with and without spin–orbit coupling (SOC), shown in Fig. SS6(j) for Te-I, reveal an indirect electronic band gap located at the high-symmetry point H. The computed band gaps are 0.49 (0.30) eV with (without) SOC, in good agreement with the experimental value of 0.33 eV,2. These results classify Te-I as a narrow-gap semiconductor. Within the energy range from 1-1 eV to 0 eV, the valence band at the H point is fourfold degenerate when SOC is neglected. The inclusion of SOC lifts this degeneracy, resulting in two non-degenerate states and one doubly degenerate state at lower energy. The region between the high-symmetry points L–H–A highlights features of the conduction band where doubly degenerate states exhibit linear dispersion—indicative of Weyl nodes located close to the Fermi level.
@@
Figs. S6(a) and (b) present the electronic band structures of $\alpha$-Te and $\beta$-Te, respectively, calculated using the MLWF-HSE06 exchange–correlation functional, both with and without spin–orbit coupling (SOC). The inclusion of SOC leads to a reduction in the band gap for both phases. Specifically, for $\alpha$-Te, the band gap decreases from $1.04$ eV to $0.75$ eV upon inclusion of SOC, in excellent agreement with previous theoretical reports,78, 75, 9. Similarly, for $\beta$-Te, the band gap decreases from $1.77$ eV to $1.44$ eV with SOC, consistent with earlier theoretical studies,78, 9, 5, 44.
@@
Nevertheless, it is important to note that non-trivial topological phases can arise under suitable external conditions. Topological phase transitions may be induced by perturbations such as mechanical strain, magnetic impurities, or doping. As discussed in Ref. 50, for instance, applying isotropic strain can drive tellurene from a trivial state into a topological phase. For completeness, the $\langle S_{x}\rangle$ spin components of the: pentagonal phase are shown in Figs. 7(c,d), buckled kagome phase [Figs. 7(e,f)], and buckled square phase [Figs. 7(g,h)].
@@
Fig. 8(a) presents the band structure of the Te helicoidal nanowire (Te-h), computed using the MLWF-HSE06 functional both with and without SOC. The inclusion of SOC consistently narrows the band gap, reducing it from 2.49 eV to 2.23 eV. These values align well with previous theoretical reports,35. In both cases, the band structure displays predominantly flat dispersion. This combination of a finite band gap and quasi-flat bands makes Te-h an appealing material for photonic applications,19. Fig. 8(b) shows the inclusion of SOC lifts band degeneracies in a manner similar to that observed in Te-I along the $\Gamma$–A direction. Along the high-symmetry path of Te-h, four band crossings are identified: two occurring at the same $K$-point—P1 in the valence band at $2.7,\text{eV}$ and P2 in the conduction band near $2,\text{eV}$—and two additional crossings near the $\Gamma$-point, P3 at $2.7,\text{eV}$ and P4 around $2,\text{eV}$. Fig. 8(c) presents the PDOS for Te-h, showing that the $p_{x}$, $p_{y}$, and $p_{z}$ orbitals contribute comparably to both the VBM and CBM. However, the CBM exhibits a more pronounced contribution from the $p_{x}$ orbital. Due to quantum confinement, Te-h displays highly localized electronic states, which is reflected in the discrete features of the band structure and in the pronounced peaks observed in the DOS.
@@
The effective masses [Table S2] were extracted from the band structure near the inflection points at the VBM and CBM. Te-h effective masses are suggestive of potentially high mobility69, 4, 41. Moreover, the modest reduction in electron and hole mobility in Te-h can help mitigate current leakage in nanoelectronic devices 66. In units of the free-electron mass, the electron effective mass of Te-h is 0.484, while the hole effective mass is 0.817 (see Table S2).
In Table 2, the topological invariant $\mathbb{Z}_{2}$ for the 2D phases of tellurium is presented. The invariants are calculated following the method described in Ref.,64. The buckled kagome and buckled square phases exhibit non-trivial topology, whereas $\alpha$-Te, $\beta$-Te, and the buckled pentagonal phase are topologically trivial. These topological classifications are consistent with the spin textures shown in Fig. 7.
Using experimentally reported lattice parameters40, the planar hexagonal tellurene phase is found to exhibit semimetallic behavior and a nontrivial $\mathbb{Z}_{2}=1$ index when SOC is included, consistent with previous experimental observations. However, our phonon calculations indicate that the free-standing planar hexagonal lattice is dynamically unstable. We therefore investigated stabilization mechanisms through strain engineering and surface functionalization. Under 5% isotropic in-plane strain, as well as under one-side hydrogen passivation, the system undergoes a transition to a gapped electronic state while preserving a nontrivial $\mathbb{Z}_{2}=1$ invariant. In these stabilized configurations, the SOC-induced gap and the winding of the Wilson loop unambiguously identify a quantum spin Hall phase.
@@
For the 2D phases, the effective masses for electrons (holes) in $\alpha$-Te are 0.108 (0.135). In $\beta$-Te, the electron effective masses are 1.009 (along X) and 0.203 (along Y), while the hole effective masses are 0.368 (X) and 0.127 (Y), indicating strong transport anisotropy. Our results suggest that both $\alpha$-Te and $\beta$-Te could have higher electron and hole mobilities than structurally or symmetrically similar materials such as 2H-MoS2 6, 67 and phosphorene 57, 34. While $\alpha$-Te exhibits higher mobilities for both charge carriers overall, $\beta$-Te shows pronounced anisotropy due to its geometry, with significantly reduced mobility along the armchair direction ($\Gamma$–X) compared to the zigzag direction ($\Gamma$–Y).